In a university class that includes 120 students, mostly non-traditional, the Mean age is 32, with a standard deviation 12 students, and with a narmal distribution. Six of these students are international students. What is the probability that, if one student is picked at a random, he/she be a teenager (12 to 19 years old)? You need to (somewhat) critically think about this before tackling it!! 1. 2. What is the probability that if ONE student is selected at random, he/she will be án international student? 3. What is the probability that if ONE student is selected at random, his/her age be between 28 to 30 years old? 4. What is the probability thatif on student is selected at random, his/her age be batween 28 to 36 years old? Suppose we select, at random, a sample of 9 students from this class. What is the probability that the MEAN age of that sample will be more than 38 years old? 5. Show your work! Write the relevant Formula and the graph!
Answer #1

P(international studnet)=6/120

P(28<x<30)=P(28-32/12<z<30-32/12)=P(-0.33<z<-0.167)=0.1967

P(28<x<36)=P(-0.33<z<0.33)=0.2586

P(xbar>38)=P(z>38-32/(12/sqrt(9))=P(z>1.5)=0.0668